I solve this by using elimination and show my work.

4x-3y=-1
x+2y=19
I need help plz. I have been trying to solve this for three days now.

Answer by Jstrasner(112)

Hey,
Instead of elimination, why dont you try subsitution:
isolate the x in the second equation:
x = 19 - 2y
Then substitute this x into the first equation:
4(19-2y) - 3y = -1
76 - 8y -3y = -1
Then we distribute the variables:
77 = 11y
y = 7
Then we substitute this y into either equation:
x + 2(7) = 19
x + 14 = 19
x = 5
I hope this helps!